Suppose that there are only two goods produced in the economy, Call center services and banking services. Prices (P), quantities (Q) and the number of workers (W) occupied in the production of each good for year 1 and year 2 are given by the following table Year 1 Year 2 Two products P1 W1 P2 Q2 W2 Call centre 10 100 50 12 100 50 Banking 10 200 50 12 230 60 b. Using year 1 prices, what is real GDP in year 2? What is the growth rate of real GDP? What is labour productivity growth between year 1 and year 2 for the whole economy? Now suppose that banking services in year 2 are not the same as banking services in year 1 because they include internet banking, which year 1 banking services didn’t include. The technology for internet banking was available in year 1 but the price of banking services with internet banking in year 1 was $13 and no one chose that package. However, in year 2 the price of banking services with internet banking was $12 and everyone chose to have that package in year 2


1
Expert's answer
2021-11-30T17:43:14-0500

Solution:

a.). Nominal GDP is derived by multiplying the current year's quantity output by the current market price.

Call center services:

Year 1 = 10 "\times" 100 = 1,000

Banking Services:

Year 1 = 10 "\times" 200 = 2,000

Nominal GDP Year 1 = 1,000 + 2,000 = 3,000

 

Call center services:

Year 2 = 12 x 100 = 1,200

Banking Services:

Year 2 = 12 "\times" 230 = 2,760

Nominal GDP Year 1 = 1,200 + 2,760 = 3,960

 

b.). Real GDP is derived by using base-year prices and multiplying them by current year quantities for all the goods and services produced in an economy.

Call center services:

Year 2 = 10 "\times" 100 = 1,000

Banking Services:

Year 2 = 10 "\times" 230 = 2,300

Real GDP Year 2 = 1,000 + 2,300 = 3,300

 

The growth rate of Real GDP = (Real GDP in Year 2 – Real GDP in Year 1)/ Real GDP in Year 1

 

Growth rate of Real GDP = (3,300 – 3,000)/3000 = 300/3000 = 10%

Growth rate of Real GDP = 10%

 

Labor productivity growth Year 1 = Total output/Number of workers = 3,000/100 = 60

Labor productivity growth Year 2 = Total output/Number of workers = 3,960/110 = 30

 

In year 2 the price of banking services with internet banking was $12, which was favorable than in year 1 which was $13 and more expensive.